# Exciton crystal melting and destruction by disorder in bilayer quantum hall system with total filling factor one

### v1

@@ Abstract
Bilayer quantum hall system with total filling factor 1 was studied in the regime of heavy layer imbalance in a recent transport experiment [1], with intriguing new findings. We demonstrate in this paper that 1) the exciton Wigner crystal in this regime can melt into a superfluid phase, giving rise to re-entrant superfluid behavior; 2) in the presence of disorder, electron and hole Wigner crystals in the two layers go through a locking/decoupling transition as layer separation increases, resulting in a sudden change in the counter flow conductance. Comparison will be made with the findings of Ref. [1].
@@ I Introduction
We start by briefly summarising the relevant observations and basic idea/conclusion of our theoretical work. The experimentalists pass a (drive) current through one of the layers, and measure the current and/or voltage response of the same as well as opposite layer; the latter corresponds to drag response [26]. Symmetric and antisymmetric combinations of these responses form normal and counter flow response functions; the latter is usually attributed to the flow of interlayer excitons which are bound pairs of electron in one layer and hole in the other, assuming they are present and dominate the counter flow transport channel. The excitons, on the other hand, may either condense to form a superfluid (SF), or crystallize and form an insulating Wigner crystal (WC) state. We will demonstrate that under appropriate conditions an exciton Wigner crystal may melt into a superfluid state, giving rise to re-entrant superfluid behavior in the counter flow channel seen in the experiment. We further demonstrate that presence of uncorrelated disorder potential in the two layers can disrupt the formation of the interlayer excitons, driving a transition between exciton Wigner crystal and decoupled electron and hole Wigner crystals in each layer. This transition manifests itself in some transport anomalies observed in the counterflow channel.
The rest of the paper is organized as following. In Sec. II, we calculate the critical temperature of bilayer exciton superfluid using two previously established effective models[8, 11] at layer imbalance $1-|\Delta\nu|\ll 1$, and demonstrate it is often higher than the melting temperature of exciton Wigner crystal. As a result the crystal melts into a superfluid when this is the case. In Sec. III we consider the interplay of disorder and interlayer coupling and analyse the competition between them. Clearly interlayer Coulomb coupling drives formation of interlayer excitons, while uncorrelated disorder favors formation of decoupled electron and hole Wigner crystals in each layer. By comparing the energy gains from exciton formation and uncorrelated electron and hole WC distortion in the two layers, we obtain the phase diagram of the system. Some concluding remarks are provided in Sec. IV.
@@ II.1 Phase stiffness and Kosterlitz-Thouless temperature of exciton superfluid
To begin with, we setup the notations here. Let $\nu_{1}=\nu_{\uparrow}=1-\delta,\nu_{2}=\nu_{\downarrow}=\delta$, we have $\Delta\nu=(1-2\delta)=\cos\theta=2(S_{\uparrow}-S_{\downarrow})=m^{z}$ and $\delta=\frac{1-\Delta\nu}{2}=\sin^{2}(\theta/2)$, density of electron in one layer $n=\delta/2\pi=\sin^{2}(\theta/2)/2\pi$.
@@ II.1.2 Dilute dipolar exciton
Thereafter superfluid phase stiffness $\rho_{s}=\frac{n}{m}$ can be obtained from $n\mathbf{v}_{s}=\rho_{s}\nabla\theta$ and $\mathbf{v}_{s}=\nabla\theta/m$.
@@ II.2 Melting temperature of exciton Wigner crystal and phase diagrams
Figure 1: Finite temperature phase diagrams near $\Delta\nu=1$ based on Equations 7 and 8. The green dashed line is the natural extension of the zero temperature phase boundary between exciton superfluid and Wigner crystal phases. Blue line is the superfluid KT temperature. Orange line is the melting curve of exciton Wigner crystal. (a) Case with $d>d_{c}\simeq 2$ in which the exciton Wigner crystal can melt into either superfluid or normal liquid, depending on $\Delta\nu$. (b) Case with $d<d_{c}$ where the exciton Wigner crystal can only melt into a superfluid.
@@ III.1 Disorder potential energy
For single layer Wigner crystal of electron-type interaction and dipole-type interaction we simply the shear modulus from [49]
@@ III.2 Interlayer correlation energy cost
As we demonstrated above, the system can lower the disorder potential energy by distort the electron and hole WCs in the two layers independently, compared to that of the exciton WC. Doing that, however, decouples the two layers and destroy the excitons, resulting in an increase in the interlayer Coulomb interaction energy. In this subsection we calculate this energy cost.
@@ IV Concluding Remarks
In this paper, we analysed the competition between different phases in a bilayer quantum hall system with total filling factor 1 driven by temperature and/or disorder. Our results compare favorably with a recent experiment [1]. Particularly interesting (and surprising) among our findings is that the exciton superfluid can (often) result from melting an exciton WC. This bears remarkable similarity to the observation [53] that melting of electron WC at low filling factor results in fractional quantum Hall liquids. Similar phenomena was observed very recently in systems supporting (fractional) anomalous quantum Hall states [54]. We speculate that melting of electron or hole WC in these systems resulted in the formation of fractional anomalous quantum Hall states.

### v2

@@ Abstract
Bilayer quantum hall system with total filling factor 1 was studied in the regime of heavy layer imbalance in a recent transport experiment [Zeng2023, arXiv:2306.16995], with intriguing new findings. We demonstrate in this paper that 1) the exciton Wigner crystal in this regime can melt into a superfluid phase, giving rise to re-entrant superfluid behavior; 2) in the presence of disorder, electron and hole Wigner crystals in the two layers go through a locking/decoupling transition as layer separation increases, resulting in a sudden change in the counter flow conductance. Comparison will be made with the findings of experiments.
@@ I Introduction
We start by briefly summarising the relevant observations and basic idea/conclusion of our theoretical work. The experimentalists pass a (drive) current through one of the layers, and measure the current and/or voltage response of the same as well as opposite layer; the latter corresponds to drag response [26]. Symmetric and antisymmetric combinations of these responses form normal and counter flow response functions; the latter is usually attributed to the flow of interlayer excitons which are bound pairs of electron in one layer and hole in the other, assuming they are present and dominate the counter flow transport channel. Bounding between electrons and hole results in the suppression of free charge carrier, and hence an insulating state of net in-plane charge transport. The excitons, on the other hand, may either condense to form a superfluid (SF), or crystallize and form an insulating Wigner crystal (WC) state. We will demonstrate that under appropriate conditions an exciton Wigner crystal may melt into a superfluid state, giving rise to re-entrant superfluid behavior in the counter flow channel seen in the experiment. We further demonstrate that presence of uncorrelated disorder potential in the two layers can disrupt the formation of the interlayer excitons, driving a transition between exciton Wigner crystal and decoupled electron and hole Wigner crystals in each layer. This transition manifests itself in some transport anomalies observed in the counterflow channel. It should be noted that, there could be other phase transitions, e.g., transition between decoupled fractional quantum hall phase and superfluid at $\Delta\nu=1/3$ [27]. They will compete with the Wigner crystal phase when $|\Delta\nu|$ goes away from 1.
The rest of the paper is organized as following. In Sec. II, we calculate the critical temperature of bilayer exciton superfluid using two previously established effective models [7, 10] at layer imbalance $1-|\Delta\nu|\ll 1$, and demonstrate it is often higher than the melting temperature of exciton Wigner crystal. As a result the crystal melts into a superfluid when this is the case. In Sec. III we consider the interplay of disorder and interlayer coupling and analyze the competition between them. Clearly interlayer Coulomb coupling drives formation of interlayer excitons, while uncorrelated disorder favors formation of decoupled electron and hole Wigner crystals in each layer. By comparing the energy gains from exciton formation and uncorrelated electron and hole WC distortion in the two layers, we obtain the phase diagram of the system. Some concluding remarks are provided in Sec. IV.
@@ II.1 Phase stiffness and Kosterlitz-Thouless temperature of exciton superfluid
To begin with, we setup the notations here. Let $\nu_{1}=\nu_{\uparrow}=1-\delta,\nu_{2}=\nu_{\downarrow}=\delta$, we have $\Delta\nu=(1-2\delta)=\cos\theta=2(S_{\uparrow}-S_{\downarrow})=m^{z}$ and $\delta=\frac{1-\Delta\nu}{2}=\sin^{2}(\theta/2)$, density of electron in one layer
$n=\delta/2\pi=\sin^{2}(\theta/2)/2\pi$ (1)
@@ II.1.2 Dilute dipolar exciton
Thereafter superfluid phase stiffness $\rho_{s}=\frac{n}{m}$ can be obtained from $n\mathbf{v}_{s}=\rho_{s}\nabla\theta$ and $\mathbf{v}_{s}=\nabla\theta/m$ where $n$ is given in (1).
@@ II.2 Melting temperature of exciton Wigner crystal and phase diagrams
Figure 1: Finite temperature phase diagrams near $\Delta\nu=1$ based on Equations 8 and 9. The green dashed line is the natural extension of the zero temperature phase boundary between exciton superfluid and Wigner crystal phases. Blue line is the superfluid KT temperature. Orange line is the melting curve of exciton Wigner crystal. (a) Case with $d>d_{c}\simeq 2$ in which the exciton Wigner crystal can melt into either superfluid or normal liquid, depending on $\Delta\nu$. (b) Case with $d<d_{c}$ where the exciton Wigner crystal can only melt into a superfluid. It should be noted that region far from $\Delta\nu=1$ shall not be taken too literally.
@@ III.1 Disorder potential energy
For single layer Wigner crystal of electron-type interaction and dipole-type interaction we simply take the shear modulus from [50]
@@ III.2 Interlayer correlation energy cost
As we demonstrated above, the system can lower the disorder potential energy by distorting the electron and hole WCs in the two layers independently, compared to that of the exciton WC. Doing that, however, decouples the two layers and destroy the excitons, resulting in an increase in the interlayer Coulomb interaction energy. In this subsection we calculate this energy cost.
@@ IV Concluding Remarks
In this paper, we analyzed the competition between different phases in a bilayer quantum hall system with total filling factor 1 driven by temperature and/or disorder. Our results compare favorably with a recent experiment [25]. Particularly interesting (and surprising) among our findings is that the exciton superfluid can (often) result from melting an exciton WC. This bears remarkable similarity to the observation [54] that melting of electron WC at low filling factor results in fractional quantum Hall liquids. Similar phenomena was observed very recently in systems supporting (fractional) anomalous quantum Hall states [55]. We speculate that melting of electron or hole WC in these systems resulted in the formation of fractional anomalous quantum Hall states. We also note that it is in principle possible to have the WC and SF orders coexist, resulting in an exciton supersolid. It is a very interesting future direction of research to look for such a novel phase, both experimentally and theoretically.